nora is 14 years old. belinda is 21/2 times her age. how old is belinda answers a:7 b:28 c:30 d:35
Answer by alicealc(293)

from the answers choice, I think you mean Belinda's age is 2 1/2 times Nora's age, not 21/2 times Nora's age.
So, Belinda is 2 1/2 * 14 = 5/2 * 14 = 5 * 7 = 35 years old (d)